A. A doctor writes a prescription for 20 mg of Tobramycin antibiotic saline solution. The solution is to be injected intravenously one time a day. On hand we have a 30 mg/ 2ml vial of Tobramycin, do the necessary calculations to give the patient the correct single dosage.
1
Expert's answer
2012-08-21T10:44:38-0400
We need 20mg/30mg = 2/3 of the given vial, so the single dosage need to be 2/3*2ml = 4/3 ml. We can dissolve the given vial by 4ml of saline. Then volume will be 6ml and then the dosage needed is 2/3*6 = 4ml.